OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the points (-2,-1) and (3,5).

OpenStudy (anonymous):

@TylerD

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ -1}})\quad &({\color{red}{ 3}}\quad ,&{\color{blue}{ 5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (tylerd):

\[\frac{ 5-(-1) }{ 3-(-2) }\]

OpenStudy (tylerd):

M = \[\frac{ 6 }{ 5 }\]

OpenStudy (tylerd):

\[y-(-1)=\frac{ 6 }{ 5 }(x-(-2))\] \[y+1=\frac{ 6 }{ 5 }x+\frac{ 12 }{ 5 }\] the finaly subtract 1 from both sides we get \[y=\frac{ 6 }{ 5 }x+1.4\]

OpenStudy (tylerd):

when you subtract a negative it becomes + thats why y-(-1) becomes y+1

OpenStudy (tylerd):

same for the x-(-2) it becomes x+2 2 times 6/5 = 12/5

OpenStudy (anonymous):

Write an equation of a line that is perpendicular to y = 2x + 3 and passes through (3,4).

OpenStudy (anonymous):

thk u so much Tyler just a few more n ill b done

OpenStudy (tylerd):

so perpendicular lines are lines that form a right angle (90 degrees) where they intersect

OpenStudy (tylerd):

so if our slope is M=2 a slope for the perpendicular line would be M=-0.5 so the slope is the negative reciprocal reciprocal of 2 = 1/2 make it negative -1/2

OpenStudy (anonymous):

hmmmm interesting keep goin Tyler

OpenStudy (tylerd):

y=-0.5x+3 is perpendicular to y=2x+3

OpenStudy (tylerd):

wait

OpenStudy (tylerd):

did not notice it says passes through (3/4)

OpenStudy (tylerd):

er (3,4)

OpenStudy (tylerd):

so in this case we have to solve y=-0.5x+b

OpenStudy (anonymous):

yes I c that

OpenStudy (tylerd):

4=-0.5(3)+b 4=-1.5+b b=5.5

OpenStudy (anonymous):

hmmmm interesting

OpenStudy (tylerd):

so the solution is y=-0.5x+5.5

OpenStudy (anonymous):

can u help with my math @mathkid10

OpenStudy (anonymous):

im here u just c meh

OpenStudy (anonymous):

*cant

OpenStudy (anonymous):

@TylerD pls help meh *~*

OpenStudy (anonymous):

my internet is actin up again

OpenStudy (tylerd):

close and start a new

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sure

OpenStudy (tylerd):